OpenStudy (anonymous):

Hi I am having trouble trying to figure out how to find the area of the region enclosed by one loop of the curve r=10sin(2theta)?

OpenStudy (anonymous):

\[\int\limits_{0}^{360} \int\limits_{0}^{10\sin(2\theta)} r*dr * d \theta \] because r varies from 0 to 10 sin(2theata). r dr dtheata is the standard way of doing a double int in polar form.

OpenStudy (anonymous):

wouldn't i have to square the r in the formula?

OpenStudy (anonymous):

I'm a little fuzzy on this, but I don't believe so. Because really you're just taking the integral of 1 * r dr dtheata, The r becomes square when you take the first integral. When you apply the bounds 10sin(2theata) drops in, the zero bound goes away.

OpenStudy (anonymous):

oh okokok thanks